OpenStudy (anonymous):

simplify 5x+30/x+4 x x+4/x^2+12+36

OpenStudy (anonymous):

\[\frac{ 5x+30 }{ x+4 }\times \frac{ x+4 }{ x ^{2}+12x+36 }\]

OpenStudy (anonymous):

@jojo4eva

OpenStudy (anonymous):

since u are multiplying u can cancel (x-4) to start with

OpenStudy (anonymous):